OpenStudy (anonymous):

y = 5 – x, solve for x y = x + b, solve for x y = a – x, solve for x

OpenStudy (anonymous):

\[y = 5 – x\] add \(x\) to both sides and get \[x+y=5\] then subtract \(y\) and get \(x=5-y\)

OpenStudy (anonymous):

method 2) \[y=5-x\] subtract \(5\) and get \[y-5=-x\] then change the signs of both sides and get \[5-y=x\]

OpenStudy (anonymous):

\[y = x + b\], solve for x subtract \(b\) from both sides what do you get?

OpenStudy (anonymous):

y = 5 – x +x +x --- ---- y + x = 5 -y -y --- ---- x = 5 - y

OpenStudy (anonymous):

@Ladyliz73 did you solve \(y=x+b\) for \(x\) ?

OpenStudy (anonymous):

x=b-y Not sure i understand still...

OpenStudy (jhannybean):

y = x + b y - b = x + b - b y - b = x that's hw you get x = y - b. The way you've written it is wrong because you're multiplying a -1 to the equation to somehow make the y negative. .

OpenStudy (cggurumanjunath):